# The first catalogue of candidate white dwarf-main sequence binaries in open star clusters: A new window into common envelope evolution

### v1

@@ 4.2 Spatial & Kinematic Properties
Figure 9: A 3-D spatial and 2-D kinematic $\chi^{2}$ association for two candidate WD+MS binaries in NGC 2287. NGC2287-c1 (triangle) and NGC2287-c2 (hexagon) are shown relative to high-probability cluster members in Cantat-Gaudin et al. (2020). Each source is coloured by its $\chi^{2}$ value, where the spatial $\chi^{2}$ is computed using right ascension ($\alpha$), declination ($\delta$) and parallax ($\pi$), whereas the kinematic $\chi^{2}$ is computed from proper motions ($\mu_{\alpha}$, $\mu_{\delta}$). Both of these candidates fall into the ‘Tier 1’ category for cluster membership, since they are both spatially and kinematically consistent with their host cluster.
@@
    Tier 1: Candidates are both spatially and kinematically consistent with their host cluster. 22/52 high-probability WD+MS candidates ($\sim$ 42%) fall in Tier 1. In our catalogue, Tier 1 candidates are flagged with Tier=‘1’ and are the most secure cluster members in our sample (see Table 1).
@@
    Tier 2: Candidates are kinematically consistent with, but spatially offset from the high-probability cluster members in Cantat-Gaudin et al. (2020). 28/52 high-probability WD+MS candidates ($\sim$54%) fall in Tier 2. Tier 2 candidates are flagged with Tier=‘2’ in our catalogue. We explore the possibility that the inconsistent spatial locations for many of these systems could be due to natal kicks in Section 6.4.
@@
    Tier 3: Candidates are both spatially and kinematically inconsistent with the host cluster. 2/52 high-probability WD+MS candidates ($\sim$4%) fall in Tier 3. Tier 3 candidates are flagged with Tier=‘3’ in our catalogue. While these candidates are located in the general vicinity of OCs–both spatially and kinematically—we caution that their cluster membership is less secure.
@@
footnotetext: An example subset of our candidate WD+MS binary catalogue. The full catalogue will be available in a machine readable format on Vizier upon acceptance of this manuscript.
4.3 Final Candidate Catalogue
As described previously, we photometrically identify 52 high-probability candidate WD+MS binaries in 38 OCs with our SVM, thus composing our catalogue. Table 1 highlights an example subset of systems and parameters included in this catalogue. A full list of catalogue parameters is described below.
@@ 5 Candidate Follow-Up Observations
Our catalogue contains 52 high-probability WD+MS binary candidates in 38 OCs, whose broadband photometric properties are consistent with those of SDSS WD+MS binaries identified by Rebassa-Mansergas et al. (2010). While spectroscopic follow-up of the full sample and detailed cases studies of individual objects will be presented in future works, here we present spectroscopy and high-cadence light curve analysis for a few example systems in order to provide context on their nature. Discussion of previous classifications for objects in our sample as well as other possible origins will be provided in Section 6 below.
@@ 5.1 Spectroscopic Follow-Up
Figure 10: An example set of spectra (left panels) and observed photometry (right panels) for three high-probability candidate WD+MS binaries in our catalogue. For each system, we show an observed spectrum/photometry (coloured), a best-fit Pickles (1998) M-dwarf model (black) and a residual spectrum (gray). Photometric bands for Pan-STARRS1 and 2MASS along with Balmer absorption features ($H\beta$, $H\gamma$, $H\delta$) are plotted for reference. Top: Alessi12-c1 (the bluest source in our catalogue); this observed spectrum was obtained with the Kast spectrograph at the Lick Observatory. For reference, we plot both the raw and smoothed spectrum. A large blue continuum flux excess and clear Balmer absorption features are present, indicating the presence of a WD in the system. At redder wavelengths, Alessi12-c1 nicely fits the shape of a M3V star. Hence, Alessi12-c1 is a clear WD+MS binary. Middle: Pleiades-c1 (the reddest source in our catalogue); this observed spectrum was obtained with the GMOS at the Gemini-North Telescope. While less dramatic than Alessi12-c1, Pleiades-c1 also exhibits an excess blue continuum flux and broad Balmer absorption features in its spectrum. While there is a slight infrared excess, Pleiades-c1 matches a M3V star at redder wavelengths. Bottom: Pleiades-c2; like Pleiades-c1, this observed spectrum was obtained with the GMOS at the Gemini-North Telescope. Neither Balmer absorption lines nor a blue continuum flux excess are observed, however both the spectrum and photometry match that of an M4V Pickles model star.
@@ 5.1.2 Description of Observed Properties
Detailed modelling and RV follow-up for Alessi12-c1, Pleaides-c1, Pleiades-c2, and other objects in our final sample are ongoing and will be presented in future work. We emphasize again, that these three objects were chosen as representative examples. SEDs for all 52 candidates are provided in Appendix C, where it is clear that other sources in our catalogue have similar photometric properties to each of those discussed in detail here.
@@ 5.2 Light Curve Data
Although more ambiguous than spectroscopy, observed system variability can provide context for the nature of an astronomical system. While light curves containing eclipses are the strongest variability indicators of both the presence of a binary and its period, eclipsing binaries are unfortunately rare. However, light curves exhibiting clear, short-period variability could indicate the presence of ellipsoidal modulations, which occur due to tidal distortions of stars in tight binaries. In addition, sinusoidal variability can be observed at a variety of timescales due to the rotation of M-dwarfs. When located in a close binary, they may be expected to be tidally-locked (e.g. Muirhead et al., 2013), leading to shorter periods (although rotation periods of $\lesssim$ 1 day are also observed in young, single M-dwarfs e.g. Newton et al. 2017). In practice, it is highly difficult to disentangle variability from binary ellipsoidal modulations and rapidly rotating M-dwarfs without radial velocity monitoring. However, since light curves with short-timescale variability can identify targets which may be short-period binaries (as expected for post-CE systems), we search for variability of high-probability candidates in our catalogue.
To identify possible variability, we perform an archival search for light curve data of each high-probability WD+MS candidate in the Kepler/K2 Mission (Howell et al., 2014), Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2015) and Zwicky Transient Facility (ZTF, Masci et al., 2019) databases. We query Kepler/K2 and TESS light curves using the Python package Lightkurve444 https://docs.lightkurve.org/ (Lightkurve Collaboration et al., 2018) and ZTF directly through the NASA/IPAC Infrared Science Archive API [^5] https://irsa.ipac.caltech.edu/docs/program_interface/ztf_lightcurve_api.html. In Figure 11, we present both raw and phase folded light curves for three high-probability WD+MS binary candidates (Alessi12-c1, Pleiades-c2 and Pleiades-c3). Alessi12-c1 and Pleiades-c2 were also shown in the spectroscopy section above. Pleiades-c3 is another candidate with slightly bluer colours, which is labeled with the number “4” in Figure 8.
@@
    Alessi12-c1: This light curve was obtained from the ZTF catalogue. When queried, Alessi12-c1 had 919 high-quality (catflags $<32768$) $r$-band observations spanning $\sim$4.7 years. The median time between sequential observations is $\sim$30 minutes, with $\sim$64of observations taken within $<$1 day of another. This is driven primarily by a few high cadence observing campaigns between years 2 and 3 (see Figure 11; left panel). Outside these times, the median cadence between observations is two days. The estimated period is 0.29 days.
@@ 6 Discussion
In the sections above, we developed a machine learning framework to identify candidate WD+MS binaries in Milky Way OCs based on their broadband photometry. In total, after vetting the quality of each candidate, we present a catalogue of $\sim$50 high-probability candidates. Here, we discuss our catalogue in the context of the three previously known WD+MS binaries in OCs (Section 6.1), previous classifications for a subset of our sample (Section 6.2), other possible origins for objects in our catalogue (Section 6.3), and implications of the kinematics observed in our sample for natal kicks of WD+MS binaries (Section 6.4).
@@ 6.1 Other Known WD+MS Systems in OCs
As described in Section 1, there are currently only three other WD+MS binaries known in OCs. V471 Tau (Nelson & Young, 1970; Young & Capps, 1971; Muirhead et al., 2022) and HZ9 (Stauffer, 1987) are the only two post-CE WD+MS binaries that have been associated with an OC. Additionally, Richer et al. (2019) present a magnetic WD+MS binary in NGC 2422. None of these three objects are contained within our final catalogue. Here, we discuss why they were not identified.
@@
The magnetic WD+MS cluster binary identified in NGC 2422 by Richer et al. (2019) is an excellent candidate to probe CE evolution. While an orbital period is not yet measured for this system, spectroscopy indicates an obvious WD and its kinematics show that it is conclusively associated with an OC. NGC 2422 is examined in this study, however the apparent Gaia magnitude of this binary $m_{G}=19.8$ is below our magnitude threshold of $m_{G}<19.5$. Hence, this source is not detected in this catalogue, but represents a strong candidate for radial velocity follow-up observations.
@@ 6.2 Previous Classifications of Candidates
To further understand the types of systems identified in this catalogue, we query all of our high-probability candidates in the SIMBAD astronomical database (Wenger et al., 2000). Of our 52 high-probability candidates, only six sources have SIMBAD classifications, where the object types are described below. We note that none of our candidates have ever been classified as WD+MS binaries. We also examine how many of the candidates selected by our method were identified as likely cluster members in the catalog of Cantat-Gaudin et al. (2020) (which we use as the kinematic baseline for our target selection in Section 2.2).
@@ 6.2.2 Eruptive Variable Stars
Pleiades-c2 and Pleiades-c3 (designated ‘V754 Tau’ and ‘V880 Tau”, respectively) are classified as eruptive variable stars in the General Catalogue of Variable Stars (GCVS; Samus’ et al., 2017). Originally identified in Haro et al. (1982), the GCVS abbreviation for these stars is ‘UV’, which represent the class of ‘UV Ceti’ type variable stars. UV Ceti stars are highly magnetic K or M-type dwarfs who undergo flaring on the order of seconds to minutes (e.g. Lacy et al., 1976).
@@ 6.2.5 Open Cluster Members
Three high-probability candidates in the Pleiades (Pleiades-c1, Pleaides-c2 and Pleiades-c3) are listed as cluster members in the Cantat-Gaudin et al. (2020) catalogue. Each source is indicated to be a cluster member with a probability of $P=1.0$, which is consistent with the low spatial and kinematic $\chi^{2}$ values for these candidates measured in Section 4.2. NGC6568-c1 is also present in the Cantat-Gaudin et al. (2020) catalogue, with a cluster probability of $P=0.2$ (likely because NGC6568-c1 exhibits kinematics that are off cluster centre). These are the only high-probability WD+MS candidates in our catalogue that are identified as cluster members in Cantat-Gaudin et al. (2020), however this is likely due to the fact that (i) all cluster members used in the Cantat-Gaudin et al. (2020) catalogue have an apparent Gaia DR2 magnitude of $m_{G}<18$ (whereas 48 of 52 high-probability candidates are fainter than this) and (ii) the large spatial cluster offsets for many of our systems could reduce membership probabilities.
@@ 6.3 Other Possible Candidate Origins
From our spectra in Figure 10, it is clear that WD+MS binaries are present in our catalogue. Since spectroscopy would be required to determine the nature of all systems, we investigate a variety of other possible astronomical sources that could be present in our sample. Each possible object is summarized below, where the respective photometric properties of each population are presented in a Gaia CMD and four colour-colour spaces in Figure 12.
Figure 12: Colour-colour and colour-magnitude spaces of a variety of astrophysical sources with similar photometry to SDSS WD+MS binaries in Rebassa-Mansergas et al. (2010). Here, we show rapidly rotating M-dwarfs (red triangles) from Kado-Fong et al. (2016), cataclysmic variables (gold hexagons) from Inight et al. (2023) and single WDs (blue diamonds) from Gentile Fusillo et al. (2021) relative to our high-probability candidate WD+MS binaries (orange stars) and our training samples (SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010) and a set of MS cluster members from Cantat-Gaudin et al. (2020) as purple and gray circles, respectively). We conclude that our catalogue is unlikely to have any significant contamination from single WDs, but our sample may contain CVs (in addition to detached WD+MS binaries). Some contamination from young, active M-dwarfs is possible for systems in our catalogue that lie close the the ZAMS.
@@ 6.3.1 Cataclysmic Variables
To investigate how CVs are classified with our SVM, we utilize the Inight et al. (2023) sample of CVs observed in SDSS I to IV. Inight et al. (2023) identify 507 CVs, representing the largest sample of spectroscopically confirmed CVs in the optical regime. After cross-matching this catalogue with 2MASS, Gaia DR3 and Pan-STARRS1 and ensuring robust photometry (Section 2.2.2 and 2.2.3), 92 CVs remain in our final sample.
From Figure 12, we see that despite being generally bluer and brighter, there is significant overlap with the Rebassa-Mansergas et al. (2010) WD+MS binaries in a Gaia CMD. As expected, the colours of CVs and WD+MS binaries are highly similar, where the CVs are mostly indistinguishable from the WD+MS binaries. These photometric similarities lead to $56\%$ of the CV sample being classified as high-probability WD+MS candidates. While the majority of our WD+MS candidates are redder than the CVs in a Gaia CMD, follow-up spectroscopy would be beneficial to distinguish between detached and accreting WD+MS binaries.

### v2

@@ 4.2 Spatial & Kinematic Properties
Figure 9: A 3-D spatial and 2-D kinematic $\chi^{2}$ association for two candidate WD+MS binaries in NGC 2287. NGC2287-c1 (triangle) and NGC2287-c2 (hexagon) are shown relative to high-probability cluster members in Cantat-Gaudin et al. (2020). Each source is colored by its $\chi^{2}$ value, where the spatial $\chi^{2}$ is computed using right ascension ($\alpha$), declination ($\delta$) and parallax ($\pi$), whereas the kinematic $\chi^{2}$ is computed from proper motions ($\mu_{\alpha}$, $\mu_{\delta}$). Both of these candidates fall into the ‘Tier 1’ category for cluster membership, since they are both spatially and kinematically consistent with their host cluster.
@@
    Tier 1: Candidates are both spatially and kinematically consistent with their host cluster. 22/52 high-probability WD+MS candidates ($\sim$ 42%) fall in Tier 1. In our catalog, Tier 1 candidates are flagged with Tier=‘1’ and are the most secure cluster members in our sample (see Table 1).
@@
    Tier 2: Candidates are kinematically consistent with, but spatially offset from the high-probability cluster members in Cantat-Gaudin et al. (2020). 28/52 high-probability WD+MS candidates ($\sim$54%) fall in Tier 2. Tier 2 candidates are flagged with Tier=‘2’ in our catalog. We explore the possibility that the inconsistent spatial locations for many of these systems could be due to natal kicks in Section 6.4.
@@
    Tier 3: Candidates are both spatially and kinematically inconsistent with the host cluster. 2/52 high-probability WD+MS candidates ($\sim$4%) fall in Tier 3. Tier 3 candidates are flagged with Tier=‘3’ in our catalog. While these candidates are located in the general vicinity of OCs–both spatially and kinematically—we caution that their cluster membership is less secure.
@@
footnotetext: This table represents an example subset of our candidate WD+MS binary catalog.
4.3 Final Candidate catalog
As described previously, we photometrically identify 52 high-probability candidate WD+MS binaries in 38 OCs with our SVM, thus composing our catalog. Table 1 highlights an example subset of systems and parameters included in this catalog. A full list of catalog parameters is described below.
@@ 5 Candidate Follow-Up Observations
Our catalog contains 52 high-probability WD+MS binary candidates in 38 OCs, whose broadband photometric properties are consistent with those of SDSS WD+MS binaries identified by Rebassa-Mansergas et al. (2010). While spectroscopic follow-up of the full sample and detailed cases studies of individual objects will be presented in future works, here we present spectroscopy and high-cadence light curve analysis for a few example systems in order to provide context on their nature. Discussion of previous classifications for objects in our sample as well as other possible origins will be provided in Section 6 below.
@@ 5.1 Spectroscopic Follow-Up
Figure 10: An example set of spectra (left panels) and observed photometry (right panels) for three high-probability candidate WD+MS binaries in our catalog. For each system, we show an observed spectrum/photometry (colored), a best-fit Pickles (1998) M-dwarf model (black) and a residual spectrum (gray). Photometric bands for Pan-STARRS1 and 2MASS along with Balmer absorption features ($H\beta$, $H\gamma$, $H\delta$) are plotted for reference. Top: Alessi12-c1 (the bluest source in our catalog); this observed spectrum was obtained with the Kast spectrograph at the Lick Observatory. For reference, we plot both the raw and smoothed spectrum. A large blue continuum flux excess and clear Balmer absorption features are present, indicating the presence of a WD in the system. At redder wavelengths, Alessi12-c1 nicely fits the shape of a M3V star. Hence, Alessi12-c1 is a clear WD+MS binary. Middle: Pleiades-c1 (the reddest source in our catalog); this observed spectrum was obtained with the GMOS at the Gemini-North Telescope. While less dramatic than Alessi12-c1, Pleiades-c1 also exhibits an excess blue continuum flux and broad Balmer absorption features in its spectrum. While there is a slight infrared excess, Pleiades-c1 matches a M3V star at redder wavelengths. Bottom: Pleiades-c2; like Pleiades-c1, this observed spectrum was obtained with the GMOS at the Gemini-North Telescope. Neither Balmer absorption lines nor a blue continuum flux excess are observed, however both the spectrum and photometry match that of an M4V Pickles model star.
@@ 5.1.2 Description of Observed Properties
Detailed modelling and RV follow-up for Alessi12-c1, Pleaides-c1, Pleiades-c2, and other objects in our final sample are ongoing and will be presented in future work. We emphasize again, that these three objects were chosen as representative examples. SEDs for all 52 candidates are provided in Appendix C, where it is clear that other sources in our catalog have similar photometric properties to each of those discussed in detail here.
@@ 5.2 Light Curve Data
Although more ambiguous than spectroscopy, observed system variability can provide context for the nature of an astronomical system. While light curves containing eclipses are the strongest variability indicators of both the presence of a binary and its period, eclipsing binaries are unfortunately rare. However, light curves exhibiting clear, short-period variability could indicate the presence of ellipsoidal modulations, which occur due to tidal distortions of stars in tight binaries. In addition, sinusoidal variability can be observed at a variety of timescales due to the rotation of M-dwarfs. When located in a close binary, they may be expected to be tidally-locked (e.g. Muirhead et al., 2013), leading to shorter periods (although rotation periods of $\lesssim$ 1 day are also observed in young, single M-dwarfs e.g. Newton et al. 2017). In practice, it is highly difficult to disentangle variability from binary ellipsoidal modulations and rapidly rotating M-dwarfs without radial velocity monitoring. However, since light curves with short-timescale variability can identify targets which may be short-period binaries (as expected for post-CE systems), we search for variability of high-probability candidates in our catalog.
To identify possible variability, we perform an archival search for light curve data of each high-probability WD+MS candidate in the Kepler/K2 Mission (Howell et al., 2014), Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2015) and Zwicky Transient Facility (ZTF, Masci et al., 2019) databases. We query Kepler/K2 and TESS light curves using the Python package Lightkurve444 https://github.com/lightkurve/ (Lightkurve Collaboration et al., 2018) and ZTF directly through the NASA/IPAC Infrared Science Archive API [^5] https://irsa.ipac.caltech.edu/docs/program_interface/ztf_lightcurve_api.html. In Figure 11, we present both raw and phase folded light curves for three high-probability WD+MS binary candidates (Alessi12-c1, Pleiades-c2 and Pleiades-c3). Alessi12-c1 and Pleiades-c2 were also shown in the spectroscopy section above. Pleiades-c3 is another candidate with slightly bluer colors, which is labeled with the number “4” in Figure 8.
@@
    Alessi12-c1: This light curve was obtained from the ZTF catalog. When queried, Alessi12-c1 had 919 high-quality (catflags $<32768$) $r$-band observations spanning $\sim$4.7 years. The median time between sequential observations is $\sim$30 minutes, with $\sim$64of observations taken within $<$1 day of another. This is driven primarily by a few high cadence observing campaigns between years 2 and 3 (see Figure 11; left panel). Outside these times, the median cadence between observations is two days. The estimated period is 0.29 days.
@@ 6 Discussion
In the sections above, we developed a machine learning framework to identify candidate WD+MS binaries in Milky Way OCs based on their broadband photometry. In total, after vetting the quality of each candidate, we present a catalog of $\sim$50 high-probability candidates. Here, we discuss our catalog in the context of the three previously known WD+MS binaries in OCs (Section 6.1), previous classifications for a subset of our sample (Section 6.2), other possible origins for objects in our catalog (Section 6.3), and implications of the kinematics observed in our sample for natal kicks of WD+MS binaries (Section 6.4).
@@ 6.1 Other Known WD+MS Systems in OCs
As described in Section 1, there are currently only three other WD+MS binaries known in OCs. V471 Tau (Nelson & Young, 1970; Young & Capps, 1971; Muirhead et al., 2022) and HZ9 (Stauffer, 1987) are the only two post-CE WD+MS binaries that have been associated with an OC. Additionally, Richer et al. (2019) present a magnetic WD+MS binary in NGC 2422. None of these three objects are contained within our final catalog. Here, we discuss why they were not identified.
@@
The magnetic WD+MS cluster binary identified in NGC 2422 by Richer et al. (2019) is an excellent candidate to probe CE evolution. While an orbital period is not yet measured for this system, spectroscopy indicates an obvious WD and its kinematics show that it is conclusively associated with an OC. NGC 2422 is examined in this study, however the apparent Gaia magnitude of this binary $m_{G}=19.8$ is below our magnitude threshold of $m_{G}<19.5$. Hence, this source is not detected in this catalog, but represents a strong candidate for radial velocity follow-up observations.
@@ 6.2 Previous Classifications of Candidates
To further understand the types of systems identified in this catalog, we query all of our high-probability candidates in the SIMBAD astronomical database (Wenger et al., 2000). Of our 52 high-probability candidates, only six sources have SIMBAD classifications, where the object types are described below. We note that none of our candidates have ever been classified as WD+MS binaries. We also examine how many of the candidates selected by our method were identified as likely cluster members in the catalog of Cantat-Gaudin et al. (2020) (which we use as the kinematic baseline for our target selection in Section 2.2).
@@ 6.2.2 Eruptive Variable Stars
Pleiades-c2 and Pleiades-c3 (designated ‘V754 Tau’ and ‘V880 Tau”, respectively) are classified as eruptive variable stars in the General catalog of Variable Stars (GCVS; Samus’ et al., 2017). Originally identified in Haro et al. (1982), the GCVS abbreviation for these stars is ‘UV’, which represent the class of ‘UV Ceti’ type variable stars. UV Ceti stars are highly magnetic K or M-type dwarfs who undergo flaring on the order of seconds to minutes (e.g. Lacy et al., 1976).
@@ 6.2.5 Open Cluster Members
Three high-probability candidates in the Pleiades (Pleiades-c1, Pleaides-c2 and Pleiades-c3) are listed as cluster members in the Cantat-Gaudin et al. (2020) catalog. Each source is indicated to be a cluster member with a probability of $P=1.0$, which is consistent with the low spatial and kinematic $\chi^{2}$ values for these candidates measured in Section 4.2. NGC6568-c1 is also present in the Cantat-Gaudin et al. (2020) catalog, with a cluster probability of $P=0.2$ (likely because NGC6568-c1 exhibits kinematics that are off cluster centre). These are the only high-probability WD+MS candidates in our catalog that are identified as cluster members in Cantat-Gaudin et al. (2020), however this is likely due to the fact that (i) all cluster members used in the Cantat-Gaudin et al. (2020) catalog have an apparent Gaia DR2 magnitude of $m_{G}<18$ (whereas 48 of 52 high-probability candidates are fainter than this) and (ii) the large spatial cluster offsets for many of our systems could reduce membership probabilities.
@@ 6.3 Other Possible Candidate Origins
From our spectra in Figure 10, it is clear that WD+MS binaries are present in our catalog. Since spectroscopy would be required to determine the nature of all systems, we investigate a variety of other possible astronomical sources that could be present in our sample. Each possible object is summarized below, where the respective photometric properties of each population are presented in a Gaia CMD and four color-color spaces in Figure 12.
Figure 12: color-color and color-magnitude spaces of a variety of astrophysical sources with similar photometry to SDSS WD+MS binaries in Rebassa-Mansergas et al. (2010). Here, we show rapidly rotating M-dwarfs (red triangles) from Kado-Fong et al. (2016), cataclysmic variables (gold hexagons) from Inight et al. (2023) and single WDs (blue diamonds) from Gentile Fusillo et al. (2021) relative to our high-probability candidate WD+MS binaries (orange stars) and our training samples (SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010) and a set of MS cluster members from Cantat-Gaudin et al. (2020) as purple and gray circles, respectively). We conclude that our catalog is unlikely to have any significant contamination from single WDs, but our sample may contain CVs (in addition to detached WD+MS binaries). Some contamination from young, active M-dwarfs is possible for systems in our catalog that lie close the the ZAMS.
@@ 6.3.1 Cataclysmic Variables
To investigate how CVs are classified with our SVM, we utilize the Inight et al. (2023) sample of CVs observed in SDSS I to IV. Inight et al. (2023) identify 507 CVs, representing the largest sample of spectroscopically confirmed CVs in the optical regime. After cross-matching this catalog with 2MASS, Gaia DR3 and Pan-STARRS1 and ensuring robust photometry (Section 2.2.2 and 2.2.3), 92 CVs remain in our final sample.
From Figure 12, we see that despite being generally bluer and brighter, there is significant overlap with the Rebassa-Mansergas et al. (2010) WD+MS binaries in a Gaia CMD. As expected, the colors of CVs and WD+MS binaries are highly similar, where the CVs are mostly indistinguishable from the WD+MS binaries. These photometric similarities lead to $56\%$ of the CV sample being classified as high-probability WD+MS candidates. While the majority of our WD+MS candidates are redder than the CVs in a Gaia CMD, follow-up spectroscopy would be beneficial to distinguish between detached and accreting WD+MS binaries.